# Approximate Polymorphisms of Predicates

### v1

@@ 9 Open questions
    Many predicates satisfy a similar but more restricted type of guarantee: every generalized polymorphism $(g_{1},\dots,g_{m})$ is such that either (i) all functions are of the form $x_{i}$ or $1-x_{i}$ (for the same $i$), or (ii) some of the functions are constant, and these constants constitute a “certificate” for the predicate. This case can be handled just as the case of $P_{\land}$. A characterizations of this type of behavior can be found in [Fil25].
@@
    Determine the optimal relation between $\epsilon$ and $\delta$ in Theorem 1.5. We conjecture that the optimal dependence is polynomial or even linear. This holds in the case of the predicates $P_{m,b}$ of Theorem 1.6, and also for many functional predicates (see [CFM+22, Appendix D] for an illustrative example).

### v2

@@ 9 Open questions
    Many predicates satisfy a similar but more restricted type of guarantee: every generalized polymorphism $(g_{1},\dots,g_{m})$ is such that either (i) all functions are of the form $x_{i}$ or $1-x_{i}$ (for the same $i$), or (ii) some of the functions are constant, and these constants constitute a “certificate” for the predicate. This case can be handled just as the case of $P_{\land}$. A characterizations of this type of behavior can be found in [Filmus2025].
@@
    Determine the optimal relation between $\epsilon$ and $\delta$ in Theorem 1.6. We conjecture that the optimal dependence is polynomial or even linear. This holds in the case of the predicates $P_{m,b}$ of Theorem 1.7, and also for many functional predicates (see [CFMMS2022, Appendix D] for an illustrative example).